{x-5}=10
if x-5=10
then x=10-5
so x=5;

Found 3 solutions by lainey, Jillian101, raghavrnr:
Answer by lainey(7)

{x-5}=10.


--Some teachers like to teach this different ways, but here is the way that I learned it. I was taught that on these types of problems, you want to get the variable (x) alone.


1.) Add 5 to each side.
x-5+5= 10 + 5


2.) The two fives would cancel out, because they add up to zero.


3.) On the right side of the equation, all you have left is x. On the left side once you solve (10 + 5), which is 15, you have x=15.


4.) The solution would be x=15.


5.) To check your work you can go back and add 10 + 5. As you can see, 10 + 5 is 15, so the solution is correct.


Answer by Jillian101(1)
Answer by raghavrnr(1)

if x-5=10
then x=10-5
so x=5;